b) Continue to assume that x = 5/4, but the market for long-distance telephone service remains perfectly competitive following the merger. For what values of s will the DOJ approve this merger? (15)
1
Expert's answer
2020-06-03T10:29:54-0400
"Q=a-bp+s"


"Q=x[K+\frac{1}{2}L]"


"x[K+\frac{1}{2}L]=a-bp+s"


"s=x[K+\frac{1}{2}L]-a+bp"

if "x=\frac{5}{4}"



"s=\frac{5}{4}[K+\frac{1}{2}L]-a+bp"